OpenStudy (anonymous):

Find the axis of symmetry and the coordinates of the vertex of the graph of the function y=-4x(squared)+3

OpenStudy (anonymous):

The axis of symmetry is found at x = -b/2a In your example a = -4 b = 0 because there are no x terms. So.. x = 0/-8 = 0 x = 0 is the line of symmetry Then plug in x = 0 and that will give you the y value of the vertex. -4*0^2 + 3 = 3 So your vertex is at (0,3)

OpenStudy (anonymous):

Find the axis of symmetry and the coordinates of the vertex of the graph of the function y=1/4x(squared)+8x